how to prepare a 5moldm-3 HCl solution from 12moldm-3 and 2moldm-3 HCl solution


1
Expert's answer
2021-06-04T15:48:01-0400

v=m/M

v(HCl)=5mol

5mol=m/36.6g/mol

m=183g


v(HCl)2=12+2=14mol

m(HCl)=14mol*36.6g/mol=512.4g

w=183g/512.4g*100%=35.71%